# SurgLaVi: Large-Scale Hierarchical Dataset for Surgical Vision-Language Representation Learning

### v1

@@ 5.2 Linear Probing
Qualitative analysis. Figure 9 shows comparative qualitative phase recognition results for representative procedures. Overall, SurgCLIP produces more temporally consistent predictions and more accurate phase boundaries compared to existing models. Similar improvements are also observed with LEMON, which benefits from large-scale self-supervised pre-training. In contrast, PeskaVLP and MedSigLIP produce noisier outputs with frequent phase switches, resulting in inconsistent predictions over time.
@@
Scaling number of frames. We study the effect of varying the temporal window size (1, 8, 16, 32 frames) across all benchmarks (Fig. 10). Performance generally improves with more frames, highlighting the importance of temporal context in workflow recognition. The only exception is GraSP step recognition, where performance peaks at 8 frames, which is consistent with the fine-grained, short-duration nature of this task compared to phase recognition. With single-frame inputs, LEMON and MedSigLIP sometimes perform better than SurgCLIP, likely due to LEMON’s frame-level self-supervised training and MedSigLIP’s higher image input resolution. However, once temporal windows are introduced, SurgCLIP exhibits superior gains, demonstrating its ability to learn robust spatio-temporal representations.
@@
Scaling number of layers. We further examine how classifier depth interacts with embeddings built from different temporal windows (Fig. 11). For single frame embeddings, performance generally increases as layers are added (except for GraSP Steps, where it declines after one layer) suggesting that these representations are not fully linearly separable and benefit modestly from additional nonlinear capacity. In contrast, embeddings constructed from a 32-frame temporal window improve only up to two layers (three for AutoLaparo) before plateauing or decreasing, consistent with the idea that temporal aggregation around the frame of interest produces more linearly separable representations, where deeper classifiers provide little benefit and may even harm performance.
@@ 5.3 Temporal probing with MS-TCN
We further evaluate our learned features in a temporal probing setting using a Multi-Stage Temporal Convolutional Network (MS-TCN) head [29]), which models long-range dependencies across full videos. This experiment complements linear probing by testing whether frozen backbone features can support a temporal head that captures procedure-level dynamics. We conduct this analysis in the full-shot regime across the three surgical phase recognition benchmarks, training the MS-TCN on frame-level features extracted from each backbone.
@@
Results. Table 3 shows that SurgCLIP achieves the highest performance on Cholec80, AutoLaparo, and GraSP phases, confirming that its representations transfer effectively to temporal models. On Cholec80 and AutoLaparo, MS-TCN probing provides additional gains compared to 32-frame linear probing, highlighting the value of extended temporal context for these datasets. In contrast, on GraSP phase recognition, temporal probing does not surpass linear probing. This can be attributed to two dataset-specific factors: (1) GraSP contains long idle or non-informative intervals between annotated phases, which weakens the temporal consistency in video squences, and (2) the limited number of full-length videos reduces the effective sample size for video-level training, whereas frame-wise linear probing mitigates this constraint. Comparing variants, SurgCLIP pre-trained on SurgLaVi- performs competitively and even surpasses the model pre-trained on full dataset on the GraSP phase task. We potentially attribute this to the higher proportion of prostatectomy procedures in SurgLaVi- (Fig. 6), which improves transferability to the GraSP domain. Finally, we observe that CLIP and MedSigLIP have exactly the same performance on GraSP, both resulting in very low F1-scores. This suggests that their features provide insufficient signal for video-level workflow modeling, leading the MS-TCN to predict nearly constant class distributions.
Model | Cholec80 | AutoLaparo | GraSP | |||||
Acc | F1 | Acc | F1 | Acc | F1 | |||
CLIP | 77.24 | 61.95 | 23.95 | 11.52 | 32.22 | 4.40 | ||
MedSigLIP | 80.88 | 69.09 | 25.01 | 5.68 | 32.22 | 4.40 | ||
PeskaVLP | 85.71 | 78.07 | 73.79 | 52.98 | 17.85 | 7.61 | ||
LEMON | 85.47 | 75.77 | 81.50 | 58.57 | 57.97 | 37.83 | ||
SurgCLIP(β) | 89.98 | 82.41 | 86.57 | 72.30 | 66.89 | 49.12 | ||
SurgCLIP | 89.87 | 82.81 | 86.58 | 73.33 | 63.70 | 44.90 |
@@ 5.4 Ablation studies
To evaluate the contribution of different components in our pre-training setup, we conduct ablation studies on the temporal granularity of pre-training data and our data curation pipeline. For computational efficiency, all ablations are performed by pre-training the model for 20 epochs. We assess zero-shot performance on phase recognition in Cholec80, AutoLaparo, and GraSP, as well as step recognition in GraSP, enabling us to validate our contributions across datasets with varying surgical workflows and temporal structures.
T | S | P | Cholec80 | AutoLaparo |
x | x | ✓ | 32.25 | 35.69 | 20.00 | 10.21 | ||||
x | ✓ | x | 40.13 | 35.51 | 24.21 | 16.60 | ||||
✓ | x | x | 40.49 | 31.32 | 29.70 | 15.71 | ||||
✓ | ✓ | ✓ | 45.36 | 35.71 | 28.12 | 18.93 |
Ablation on hierarchical pre-training levels. As shown in Table 4, joint multi-level pre-training (task + step + phase) achieves the best overall zero-shot performance across datasets, confirming that integrating multiple temporal granularities provides complementary supervision. When trained on individual levels, task-only and phase-only models generally outperform phase-only pre-training. We attribute this to two factors: (1) phase-level clips are longer in duration, resulting in fewer training samples compared to finer granularities, and (2) phase captions are considerably more complex, making them harder to learn in isolation. In contrast, finer-grained levels (task, step) increase the number of available samples and provide simpler, more localized procedural descriptions, which may be easier for the model to align with visual patterns compared to phase-level samples. We hypothesize that when all levels are used jointly, the model benefits from the abundance and simplicity of fine-grained supervision and the richer procedural semantics of phase-level descriptions, enabling it to learn robust representations across multiple temporal scales and semantic granularities.
MLSC | DMF | CE | Cholec80 | AutoLaparo |
x | x | x | 36.97 | 31.62 | 23.91 | 14.50 | ||||
✓ | x | x | 38.57 | 31.74 | 27.52 | 15.71 | ||||
✓ | ✓ | x | 44.39 | 34.46 | 29.89 | 18.56 | ||||
✓ | ✓ | ✓ | 45.36 | 35.71 | 28.12 | 18.93 |
Ablation on data curation pipeline components. Table 5 reports the effect of progressively adding components of our curation pipeline. Using only sentence-level transcriber timestamps to define clip–caption pairs results in the weakest performance, as boundaries often cut across incomplete semantic units. Incorporating Multi-Level Semantic Captioning (MLSC) produces richer pairs with varied temporal spans and caption complexities, leading to clear gains. Adding Dual-Modal Filtering (DMF) further improves results by discarding non-surgical or weakly descriptive pairs, thereby reducing noise in vision–language alignment. As shown in Figure 2, this step reduces the dataset size by 35%, revealing an important insight: increasing scale alone does not guarantee better performance, data quality plays a critical role as well. This observation aligns with findings reported by Wang et al. [30]. Finally, Contextual Enrichment (CE) enhances captions with procedural context and achieves the best overall scores, with the exception of GraSP phases where DMF alone is optimal. These results show that each stage contributes complementary benefits, and together they achieve the strongest supervision for representation learning.

### v2

@@ 5.2.2 Context Optimization (CoOp)
Qualitative analysis. Figure 8 shows comparative qualitative phase recognition results for representative procedures under full-shot linear probing. Overall, SurgCLIP produces more temporally consistent predictions and more accurate phase boundaries compared to existing models. Similar improvements are also observed with LEMON, which benefits from large-scale self-supervised pre-training. In contrast, PeskaVLP and MedSigLIP produce noisier outputs with frequent phase switches, resulting in inconsistent predictions over time.
@@
Scaling number of frames. We study the effect of varying the temporal window size (1, 8, 16, 32 frames) across all benchmarks under full-shot linear probing (Fig. 11). Performance generally improves with more frames, highlighting the importance of temporal context in workflow recognition. The only exception is GraSP step recognition, where performance peaks at 8 frames, which is consistent with the fine-grained, short-duration nature of this task compared to phase recognition. With single-frame inputs, LEMON and MedSigLIP sometimes perform better than SurgCLIP, likely due to LEMON’s frame-level self-supervised training and MedSigLIP’s higher image input resolution. However, once temporal windows are introduced, SurgCLIP exhibits superior gains, demonstrating its ability to learn robust spatio-temporal representations.
@@
Scaling number of layers. We further examine how classifier depth interacts with embeddings built from different temporal windows under full-shot linear probing (Fig. 10). For single frame embeddings, performance generally increases as layers are added (except for GraSP Steps, where it declines after one layer) suggesting that these representations are not fully linearly separable and benefit modestly from additional nonlinear capacity. In contrast, embeddings constructed from a 32-frame temporal window improve only up to two layers (three for AutoLaparo) before plateauing or decreasing, consistent with the idea that temporal aggregation around the frame of interest produces more linearly separable representations, where deeper classifiers provide little benefit and may even harm performance.
@@ 5.3 Temporal probing with MS-TCN
We further evaluate our learned features in a temporal probing setting using a Multi-Stage Temporal Convolutional Network (MS-TCN) head [farha2019ms], which models long-range dependencies across full videos. This experiment complements linear probing by testing whether frozen backbone features can support a temporal head capable of capturing procedure-level dynamics. Because temporal modeling requires not only frame-level classification but also coherent and well-aligned segment boundaries, we assess both frame-level metrics (video-wise Accuracy and macro-video-wise F1) and temporal segmentation metrics (Segmental F1@), following prior work in temporal segmentation [lea2017temporal, zhang2023surgical]. We conduct this analysis in the full-shot regime across the three surgical phase recognition benchmarks. MS-TCN is trained on frame-level features extracted from each backbone, ensuring a consistent comparison across models. Additional details are provided in Section 1.4 in the Appendix.
| Metric | Model | |||||
| CLIP | MedSigLIP | PeskaVLP | LEMON | SurgCLIP(β) | SurgCLIP | |
| Cholec80 | ||||||
| Acc | 77.24±10.82 | 80.88±12.02 | 85.71±11.10 | 85.47±11.66 | 89.72±7.79 | 90.11±7.34 |
| F1 | 61.95±13.26 | 69.09±13.66 | 78.07±11.15 | 75.77±12.40 | 82.39±9.57 | 83.35±8.84 |
| F1@10 | 30.12±13.96 | 56.97±23.49 | 54.02±22.08 | 65.38±22.71 | 86.22±17.15 | 84.49±14.49 |
| F1@25 | 27.43±13.88 | 53.49±23.48 | 52.76±22.67 | 63.63±23.68 | 85.78±16.97 | 83.46±15.25 |
| F1@50 | 22.14±12.62 | 45.06±24.49 | 47.23±23.96 | 57.38±26.47 | 78.57±20.06 | 77.05±18.26 |
| AutoLaparo | ||||||
| Acc | 23.95±6.53 | 25.01±5.59 | 73.79±11.63 | 81.50±9.17 | 86.52±6.60 | 86.59±7.64 |
| F1 | 11.52±3.11 | 5.68±1.05 | 52.98±6.58 | 58.57±5.99 | 72.32±6.90 | 73.34±7.23 |
| F1@10 | 0.11±0.13 | 14.02±2.10 | 12.14±6.34 | 36.93±15.99 | 69.43±14.23 | 62.96±13.47 |
| F1@25 | 0.00±0.00 | 5.40±6.27 | 10.61±5.14 | 35.31±14.75 | 68.24±14.75 | 60.53±13.01 |
| F1@50 | 0.00±0.00 | 0.00±0.00 | 6.88±4.29 | 31.47±14.49 | 59.90±14.15 | 50.47±17.68 |
| GraSP | ||||||
| Acc | 32.22±6.52 | 32.22±6.52 | 17.85±3.02 | 57.97±4.08 | 66.58±5.97 | 63.69±6.83 |
| F1 | 4.40±6.70 | 4.40±6.70 | 7.61±11.80 | 37.83±11.06 | 48.98±13.08 | 44.89±10.28 |
| F1@10 | 0.69±8.70 | 0.69±8.70 | 1.82±0.87 | 22.79±5.68 | 38.83±6.80 | 38.55±4.20 |
| F1@25 | 0.00±0.00 | 0.00±0.00 | 0.67±0.40 | 18.08±5.42 | 32.02±6.12 | 32.10±5.55 |
| F1@50 | 0.00±0.00 | 0.00±0.00 | 0.20±0.14 | 9.05±4.66 | 18.76±5.04 | 17.83±4.67 |
@@
Results. Table 3 shows that SurgCLIP achieves the highest F1 and Accuracy performance on Cholec80 and AutoLaparo, while SurgCLIP(β) attains the best overall performance in GraSP, confirming that our learned representations transfer effectively to temporal models. On Cholec80 and AutoLaparo, MS-TCN probing provides additional gains compared to 32-frame linear probing, highlighting the value of extended temporal context for these datasets. In contrast, on GraSP phase recognition, temporal probing does not surpass linear probing. We attribute this to two dataset-specific factors: (1) GraSP contains long idle or non-informative intervals between annotated phases, which weakens the temporal consistency within videos; and (2) the limited number of full-length videos reduces the effective sample size for video-level training, whereas frame-wise linear probing mitigates this limitation.
When examining segmental F1 metrics, we observe that although SurgCLIP yields slightly higher frame-level accuracy and F1 on AutoLaparo and Cholec80, SurgCLIP(β) demonstrates substantially improved temporal consistency across nearly all thresholds. These findings suggest that the pretraining distribution in SurgLaVi- produces features with stronger temporal coherence for certain downstream tasks compared to the full SurgLaVi corpus. On GraSP phases, SurgCLIP(β) outperforms the full model across frame-level and segmental metrics, which we attribute to the higher proportion of prostatectomy procedures in SurgLaVi- (Figure 6), resulting in improved domain alignment.
Beyond our two models, the temporal probing results also reveal clear performance differences across existing baselines. CLIP provides very limited temporal signal to MS-TCN, resulting in weak, or even zero, segmental F1 scores and failing to capture procedural regularities. MedSigLIP, despite being trained on medical image-text pairs, remains insufficiently specialized for surgical dynamics and offers only modest improvements. Surgical VLMs such as PeskaVLP perform better at the frame level but still struggle with temporal consistency. Notably, LEMON performs better than CLIP, MedSigLIP, and PeskaVLP, demonstrating that large-scale self-supervised in-domain pretraining produces more informative features for temporal modeling. However, LEMON still trails behind our SurgCLIP models due to the absence of semantic procedural grounding. In contrast, SurgCLIP and SurgCLIP(β) deliver consistently superior temporal segmentation, often more than doubling the F1@10 of the next-best baseline, demonstrating that large-scale surgical hierarchical vision-language alignment is crucial for learning representations with strong temporal structure.
5.4 Upper-Bound Results: Comparison with Task-Specific Models
| Training | AutoLaparo | Cholec80 | |||||||||
| Samples | Model | Backbone | Head | Acc (%) | P (%) | R (%) | J (%) | Acc (%) | P (%) | R (%) | J (%) |
| Full | EndoNet [endonet] | CNN★ | Linear Layer★ | – | – | – | – | – | |||
| SV-RCNet [svrcnet] | CNN★ | LSTM★ | 75.6 | 64.0 | 59.7 | 47.2 | – | ||||
| TMRNet [tmrnet] | CNN★ | LSTM★ | 78.2 | 66.0 | 61.5 | 49.6 | |||||
| TeCNO [czempiel2020tecno] | CNN★ | MS-TCN★ | 77.3 | 66.9 | 64.6 | 50.7 | |||||
| Trans-SVNet [gao2021trans] | CNN★ | Transformer / TCN★ | 78.3 | 64.2 | 62.1 | 50.7 | 89.1 | 84.7 | 83.6 | 72.5 | |
| AVT [girdhar2021anticipative] | Transformer★ | Transformer★ | 77.8 | 68.0 | 82.2 | 50.7 | 86.7 | 77.3 | 82.1 | 66.4 | |
| LoViT [liu2025lovit] | Transformer★ | Transformer★ | 81.4 | 85.1 | 65.9 | 56.0 | 91.5 | 83.1 | 86.6 | 74.2 | |
| SKiT [liu2023skit] | Transformer★ | Transformer★ | 82.9 | 81.8 | 70.1 | 59.9 | 92.5 | 84.6 | 88.5 | 76.7 | |
| Surgformer [yang2024surgformer] | Transformer★ | Linear Layer★ | 85.7 | 82.3 | 75.7 | 66.7 | 92.4 | 87.9 | 89.3 | 79.9 | |
| SurgCLIP(β) | Transformer● | Linear Layer★ | 81.9 | 74.4 | 70.4 | 60.6 | 84.8 | 79.8 | 80.0 | 60.6 | |
| SurgCLIP | Transformer● | Linear Layer★ | 84.6 | 71.3 | 71.3 | 61.3 | 85.2 | 80.2 | 80.4 | 67.3 | |
| CLS Few-Shot Linear Probing (samples per class) | |||||||||||
| 1 | SurgCLIP(β) | Transformer● | Linear Layer★ | ||||||||
| 2 | |||||||||||
| 4 | |||||||||||
| 8 | |||||||||||
| 16 | |||||||||||
| 1 | SurgCLIP | Transformer● | Linear Layer★ | ||||||||
| 2 | |||||||||||
| 4 | |||||||||||
| 8 | |||||||||||
| 16 | |||||||||||
| Video Linear Probing (% of available labeled video cases) | |||||||||||
| 10% | SurgCLIP(β) | Transformer● | Linear Layer★ | ||||||||
| 50% | |||||||||||
| 10% | SurgCLIP | Transformer● | Linear Layer★ | ||||||||
| 50% | |||||||||||
| Samples | Model | Backbone | Head | F1 Phases (%) | F1 Steps (%) |
| Full | SlowFast [feichtenhofer2019slowfast] | CNN★ | Linear Layer★ | 54.7 | 37.1 |
| TAPIR [valderrama2022towards] | Transformer★ | Linear Layer★ | 58.6 | 43.8 | |
| TAPIS [ayobi2025pixel] | Transformer★ | Linear Layer★ | 63.4 | 45.8 | |
| MuST [perez2024must] | Transformer★ | Transformer★ | 68.8 | – | |
| SurgCLIP(β) | Transformer● | Linear Layer★ | 58.9 | 35.5 | |
| SurgCLIP | Transformer● | Linear Layer★ | 61.1 | 38.5 | |
| CLS Few-Shot Linear Probing (samples per class) | |||||
| 1 | SurgCLIP(β) | Transformer● | Linear Layer★ | ||
| 2 | |||||
| 4 | |||||
| 8 | |||||
| 16 | |||||
| 1 | SurgCLIP | Transformer● | Linear Layer★ | ||
| 2 | |||||
| 4 | |||||
| 8 | |||||
| 16 | |||||
| Video Linear Probing (% of available labeled video cases) | |||||
| 10% | SurgCLIP(β) | Transformer● | Linear Layer★ | ||
| 50% | |||||
| 10% | SurgCLIP | Transformer● | Linear Layer★ | ||
| 50% | |||||
Having demonstrated superior performance over existing VLMs and self-supervised foundation models in surgery across multiple evaluation settings including linear probing, temporal probing, and CoOp-based adaptation, we now benchmark SurgCLIP and SurgCLIP(β) against task-specific architectures on Cholec80, AutoLaparo, and GraSP (Phases and Steps). We position these fully-supervised methods as upper-bound performance references, as they perform end-to-end training of both backbone and heads (often with dataset-tailored architectural choices), whereas our approach relies on frozen features from our SurgLaVi-pretrained models with only a lightweight linear classification layer being tuned under various few-shot and full-shot settings described in Section 5.2. To contextualize our approach within the broader surgical phase recognition literature, we adapted our experimental protocol to align with established benchmarks for AutoLaparo and Cholec80. Specifically, we have included Accuracy, Precision, Recall, and Jaccard score used for these datasets as defined in [funke2023metrics], and employed an online inference setup to ensure fair comparison with specialized models.
Table 4 presents our results on the AutoLaparo and Cholec80 datasets. On AutoLaparo, our models trained with only 16 samples per class surpass CNN-based baselines with complex temporal heads (LSTM, TCN, Transformer, MS-TCN) across Accuracy, Precision, and Jaccard metrics, despite using only a single trainable linear layer. While performance remains below fully-trainable transformer-based architectures with sophisticated temporal aggregation methods, this gap narrows substantially when full training data is used. Under this setting, SurgCLIP achieves competitive Accuracy (84.6%), Recall (71.3%), and Jaccard score (61.3%), demonstrating strong transfer from our pretraining corpus. We observe a different trend on Cholec80, where the 16-few-shot regime proves insufficient to get close to the performance of specialized models. This reflects the larger training data disparity between AutoLaparo and Cholec80, where specialized models in Cholec80 benefit from 4 more training videos than those in AutoLaparo. However, with 50% of the training data, our models achieve performance comparable to CNN-based architectures. With full data availability, SurgCLIP matches the transformer-based AVT model and remains competitive with other heavy transformer-based models, all while requiring only a linear classification layer during training.
For GraSP phase and step recognition (Table 5), our linear probe matches CNN-based baselines using only 16 samples per class. With full training data, SurgCLIP’s linear probe outperforms TAPIR (Transformer-based) and SlowFast (CNN-based) while using minimal trainable parameters and is only 2.34% F1 lower compared to TAPIS. Performance remains below MuST, which achieves stronger results by employing substantially more parameters through multiple multi-scale transformer heads stacked on top of its transformer-based vision backbone.
Taken together, these results demonstrate that SurgCLIP delivers competitive recognition performance across diverse surgical datasets using only frozen features and a single trainable linear layer, substantially reducing training cost. Although specialized architectures that jointly optimize backbone and temporal aggregation still maintain a 7.4% advantage on average, the gap is modest relative to their significantly higher architectural complexity and training requirements.
@@ 5.5 Ablation studies
To evaluate the contribution of different components in our pre-training setup, we conduct ablation studies on the temporal granularity of pre-training data, our data curation pipeline, and our dynamic sampling during training. We assess zero-shot performance on phase recognition in Cholec80, AutoLaparo, and GraSP, as well as step recognition in GraSP, enabling us to validate our contributions across datasets with varying surgical workflows and temporal structures.
Ablation on hierarchical pre-training levels. As shown in Table 6, joint multi-level pre-training that uses fine, mid, and coarse data together achieves the best overall zero shot performance across datasets. This supports the idea that supervision from different temporal scales provides complementary benefits. When each level is used on its own, pre-training with fine-level data or mid-level data produces stronger results than pre-training with coarse-level data. In the context of traditional visual recognition of surgical workflow, visual models often perform better on coarse task such as phases because they last longer, change less frequently, and do not require the same level of temporal precision as granular tasks. However, in our multi-modal setting, fine-level and mid-level data produce a larger number of training samples, and the captions are shorter, more focused, and easier for the model to relate to the visual content, as shown in Figure 3. Coarse-level clips are longer and less numerous, and the corresponding captions are broader and more complex, which makes it more difficult for the model to learn strong associations when coarse-level data is used alone. When all levels are combined during pre-training, the model benefits from the clarity and abundance of the fine-grained supervision, while also learning from the richer procedural information contained in coarse-level descriptions.
| F | M | C | Cholec80 | AutoLaparo |
| x | x | ✓ | 32.34±6.54 | 43.70±4.95 | 20.76±4.23 | 12.85±4.03 | ||||
| x | ✓ | x | 39.72±6.42 | 43.65±3.67 | 23.73±5.44 | 13.57±3.97 | ||||
| ✓ | x | x | 39.12±7.59 | 53.48±5.62 | 31.15±7.80 | 16.81±6.42 | ||||
| ✓ | ✓ | ✓ | 50.53±8.83 | 56.37±5.54 | 34.94±7.61 | 16.53±5.44 |
| DMF | MLSC | CE | Cholec80 | AutoLaparo |
| Baseline | ||||||||||
| 35.89±5.98 | 46.10±9.31 | 24.59±7.71 | 13.35±4.17 | |||||||
| Individual components | ||||||||||
| ✓ | 35.88±6.03 | 43.42±7.26 | 31.69±7.39 | 14.33±4.43 | ||||||
| ✓ | 36.11±7.60 | 36.19±7.11 | 27.50±6.88 | 12.52±4.25 | ||||||
| ✓ | 33.42±7.11 | 24.20±5.61 | 9.55±3.14 | |||||||
| Pairwise combinations | ||||||||||
| ✓ | ✓ | 33.56±8.01 | 51.65±7.68 | 25.79±5.39 | 11.91±3.79 | |||||
| ✓ | ✓ | 42.36±8.42 | 47.34±7.56 | 32.02±8.15 | 14.57±4.51 | |||||
| Complete pipeline | ||||||||||
| ✓ | ✓ | ✓ | 50.53±8.83 | 56.37±4.54 | 34.94±7.61 | 16.53±5.44 | ||||
| Mode | Stride | Cholec80 | AutoLaparo | GraSP Phases | GraSP Steps |
| Fixed | 1 | 43.75±7.35 | 34.18±9.43 | 31.36±8.87 | 14.19±5.45 |
| 4 | 40.45±8.06 | 41.82±5.19 | 24.75±7.96 | 13.15±1.31 | |
| Dynamic | 50.53±8.83 | 56.37±4.54 | 34.94±7.61 | 16.53±5.44 | |
Ablation on data curation pipeline components. Table 7 reports the effect of individual components and their combinations in our curation pipeline. We observe that components exhibit varying effectiveness in isolation but demonstrate strong synergistic effects when combined. Examining individual components, we find that their isolated contributions are inconsistent across datasets. Dual-Modal Filtering (DMF) alone yields mixed results: it maintains baseline performance on Cholec80 while showing gains on GraSP Phases (+7.10%) but degradation on AutoLaparo (-2.68%). Multi-Level Semantic Captioning (MLSC) in isolation provides minimal improvements and even decreases performance on AutoLaparo (36.19% vs. 46.10% baseline). Most notably, Contextual Enrichment (CE) alone consistently underperforms or matches the baseline across most benchmarks, with substantial drops on Cholec80 and GraSP Steps, despite gains on AutoLaparo. This reveals that individual components, when applied in isolation, do not consistently improve performance and may even introduce noise without complementary mechanisms. However, pairwise combinations begin to reveal the synergistic nature of our pipeline. Combining DMF with MLSC achieves notable improvements over the baseline (+6.47% on Cholec80, +7.43% on GraSP Phases), demonstrating that semantic diversity (MLSC) paired with quality filtering (DMF) creates more effective training data. Interestingly, pairing DMF with CE results in stronger performance on AutoLaparo (+5.55% over baseline) but remains suboptimal on other datasets, suggesting that contextual enrichment requires semantic grounding to be fully effective. The complete pipeline with all three components achieves the best performance across all benchmarks, with substantial gains: +14.64% on Cholec80, +10.27% on AutoLaparo, +10.35% on GraSP Phases, and +3.18% on GraSP Steps compared to baseline. Notably, the full method outperforms even the best pairwise combination by significant margins (e.g., +8.17% on Cholec80 over DMF+MLSC), demonstrating that all three components are mutually reinforcing rather than independently additive for effective vision-language alignment in surgical video understanding.
Ablation on temporal sampling strategies. Table 8 compares fixed-stride sampling with our dynamic sampling strategy during pre-training. Our dynamic sampling strategy adapts the sampling stride based on segment duration while maintaining a fixed number of frames per clip: shorter, task-level segments use smaller strides for high temporal resolution, whereas longer, coarse-level segments employ larger strides to increase temporal coverage. This ensures that each clip accurately reflects the temporal coverage described by its paired caption, maintaining proper vision-language alignment while exposing the model to multiple temporal resolutions during training. The results demonstrate the effectiveness of this approach, with our dynamic sampling strategy substantially outperforming fixed-stride alternatives across all benchmarks, resulting +6.78% to +14.55% improvement over stride-1 and +9.08% to +22.19% over stride-4 across datasets.
6 Limitations
Our work shows that large scale multi-modal pretraining on hierarchically structured and automatically curated surgical video–language data can meaningfully advance generalizable representation learning in surgery, improving transferability across procedures and modalities while narrowing, although not yet closing, the gap with fully supervised and task specific models. These results highlight the promise of vision–language methods for broad surgical intelligence. Nonetheless, several limitations remain. Our pipeline works on narrated surgical videos, which are not uniformly available across all procedure types or clinical settings. In addition, narrations may not always reflect the exact moment shown on screen since surgeons often describe recent actions or upcoming steps, which can occasionally introduce temporal misalignment between captions and visual content. Furthermore, the fully automated curation process operates without expert verification, and although foundation models make it possible to process data at the scale required for effective pretraining, they may introduce noise or bias in the process. Future work could aim to leverage silent videos to complement narrated content by combining self-supervised methods with multi-modal contrastive learning to improve generalizability and robustness of learned representations.